The alarm of Clock A rings every 4 minutes, the alarm of Clock B rings every 6 minutes, the alarm of Clock C rings every 7 minutes. If the alarms of all three clock's alarms ring at 12:00nn, the next time at which at the alarms will ring at exactly the same time is?
Found 2 solutions by nerdybill, richwmiller:
Answer by nerdybill(7384)

The alarm of Clock A rings every 4 minutes, the alarm of Clock B rings every 6 minutes, the alarm of Clock C rings every 7 minutes. If the alarms of all three clock's alarms ring at 12:00nn, the next time at which at the alarms will ring at exactly the same time is?
.
Find the LCM (lowest common multiple) of 4, 6, and 7:
factor each to prime numbers:
4: 2 2
6: 2 3
7: 7
.
LCM: 2*2*3*7 = 84 minutes
.
12:00 + 84 minutes = 1:24

Answer by richwmiller(17219)

4*6*7=42*4=168 minutes at 2:48 pm
A and B will ring together every 12 minutes. 7*12=84
So all three will also ring together after 84 minutes or at 1:24 pm